OpenStudy (anonymous):

A regular octagon has side length 10.9 in. The perimeter of the octagon is 87.2 in cm and the area is 392.4 in2. A second octagon has side lengths equal to 21.8 in. Find the perimeter of the second octagon. Round to the nearest hundredth.(Points : 3)

Directrix (directrix):

Is the second octagon regular? Although the question does not specifiy that, I assume the second octagon is regular because of the statement "has side lengths equal to 21.8 in."

Directrix (directrix):

Any two regular octagons are similar. The ratio of the sides (scale factor) is the same as the ratio of the perimeters.

Directrix (directrix):

10.9 / 21.8 = 87.2 / p where p is the perimeter of the second octagon.

Directrix (directrix):

10.9 / 21.8 = 87.2 / p Cross multiply to get: 10.9 * p = (87.2)*(21.8) What do you get when you solve for p? --- @Brandon_hockinson